The following wave function |A> given as:
|A> = (9-2i) |u_1> + 4i |u_2> - |u_3 + i|u_4

(a) Find the dual vector <A|
(b) If |u_1> is under an orthonormal base, find the representation of |A> as a column vector in the basis given.
1
Expert's answer
2020-08-10T19:55:01-0400

a)


"\vec{A^*}=\frac{1}{A} \begin{pmatrix}
 9+2i & -4i & -1 & -i 
\end{pmatrix}\\\vec{A^*}=\frac{1}{\sqrt {103}} \begin{pmatrix}
 9+2i & -4i & -1 & -i 
\end{pmatrix}"

b)

"\vec{A}= \begin{pmatrix}
 9-2i \\ 4i \\ -1 \\ i 
\end{pmatrix}"